OpenStudy (anonymous):

Find the exact value of cos 330°

OpenStudy (zzr0ck3r):

11/12 of 2Pi = (11/6)pi

OpenStudy (anonymous):

cos(360-30)=cos 30=sqrt((3))/2

OpenStudy (anonymous):

look it as cos(360-30)

OpenStudy (anonymous):

i 4th quadrent cos is positive so \[\cos 30=\sqrt{3}/2\]